OpenStudy (anonymous):

the sum of three consecutive even numbers is 84. what is the smallest of these numbers

OpenStudy (anonymous):

27 28 29

OpenStudy (anonymous):

x = first number x + 1 = second number x + 2 = third number x+x+1+x+2 = 84

OpenStudy (anonymous):

so 27

OpenStudy (anonymous):

3x + 2 = 84 3x = 82 x = 27 27 is the smallest number

OpenStudy (campbell_st):

no they are even numbers so its x x+2 and x+4 so x + x+2 + x + 4 = 84 now solve for x

OpenStudy (anonymous):

Oops... campbell is right...didn't read carefully

OpenStudy (anonymous):

oh crap

OpenStudy (anonymous):

dang! this always happens!